# Disentangled Representation Learning with the Gromov-Monge Gap

### v2

@@ 3.1 From the distortion…
Recently, Uscidda and Cuturi (2023) introduced the Monge gap, a regularizer that measures whether a map $T$ transports a reference distribution at the minimal displacement cost. Building on this concept, we replace "displacement" with "distortion" to introduce the Gromov-Monge gap, a regularizer that assesses whether a map $T$ transports a reference distribution at the minimal distortion cost.
@@
The Influence of the Reference Distribution.
A crucial property of $\mathrm{DST}_{r}$ is that if $T$ transforms $r$ without distortion, it will also apply distortion-free to any distribution $s$ whose support is contained within that of $r$. Formally, if $\mathrm{DST}_{r}(T)=0$ and $s\in\mathcal{P}(\mathcal{X})$ with $\mathrm{Spt}(s)\subseteq\mathrm{Spt}(r)$, then $\mathrm{DST}_{s}(T)=0$. This raises a question for the GMG: If $T$ maps $r$ with minimal distortion, does it similarly map $s$ with minimal distortion? We answer this question positively with Prop. (3.3).
@@ 3.3 Estimation and Computation from Samples
$\begin{split}\mathrm{DST}_{{r}_{n}}(T)=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(T(\mathbf{x}_{i}),T(\mathbf{x}_{j})))^{2}\,,\end{split}$ (3)
@@
2:$\mathbf{t}_{1},\dots\mathbf{t}_{n}\leftarrow T(\mathbf{x}_{1}),\dots,T(\mathbf{x}_{n})$.
3:$\mathbf{C}_{\mathcal{X}}\leftarrow[c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})]_{1\leq i,i^{\prime}\leq n}$    $\triangleright$ usually $\mathcal{O}(n^{2}d_{\mathcal{X}})$
4:$\mathbf{C}_{\mathcal{Y}}\leftarrow[c_{\mathcal{Y}}(\mathbf{t}_{j},\mathbf{t}_{j^{\prime}})]_{1\leq j,j^{\prime}\leq n}$    $\triangleright$ usually $\mathcal{O}(n^{2}d_{\mathcal{Y}}$)
5:$\textsc{DST}\leftarrow\texttt{mean}((\mathbf{C}_{\mathcal{X}}-\mathbf{C}_{\mathcal{Y}})^{2})$triangleright $n^{2}$
6:$\varepsilon\leftarrow\varepsilon_{0}\cdot\texttt{stat}(\mathbf{C}_{\mathcal{X}})\cdot\texttt{stat}(\mathbf{C}_{\mathcal{Y}})$triangleright usually $\mathcal{O}(n^{2})$
7:$\textsc{GW}\leftarrow\texttt{GW}(\mathbf{C}_{\mathcal{X}},\mathbf{C}_{\mathcal{Y}},\varepsilon)$triangleright $\mathcal{O}(n^{3})$ or $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$
@@
Choosing $\varepsilon=0$, we recover $\mathrm{GMG}_{{r}_{n},0}(T)=\mathrm{GMG}_{{r}_{n}}(T)$. Moreover, the entropic estimator preserves positivity, as for $\varepsilon\geq 0$, $\mathrm{GMG}_{r_{n},\varepsilon}(T)\geq 0$ (see B.3). We compute $\mathrm{GW}_{\varepsilon}({r}_{n},T\sharp{r}_{n})$ using Peyré et al. (2016)’s solver introduced in 2.2. We use the implementation provided by ott-jax (Cuturi et al., 2022). In practice, we select $\varepsilon$ based on (positive) statistics from the cost matrices $\mathbf{C}_{\mathcal{X}}$,$\mathbf{C}_{\mathcal{Y}}$. We define a scale $\varepsilon_{0}$ and set $\varepsilon=\varepsilon_{0}\cdot\texttt{stat}(\mathbf{C}_{\mathcal{X}})\cdot\texttt{stat}(\mathbf{C}_{\mathcal{Y}})$. Standard options for the statistic include $\texttt{stat}\in\{\texttt{mean},\texttt{max},\texttt{std}\}$. This procedure is equivalent to running the entropic GW solver on the re-scaled cost matrices $\mathbf{C}_{\mathcal{X}}/\texttt{stat}(\mathbf{C}_{\mathcal{X}})$ and $\mathbf{C}_{\mathcal{Y}}/\texttt{stat}(\mathbf{C}_{\mathcal{Y}})$ with $\varepsilon=\varepsilon_{0}$; see App. B.2. It is a common practical trick, initially suggested for stabilizing the Sinkhorn algorithm Cuturi (2013), then extended to the entropic GW solver (Cuturi et al., 2022). Additionally, as suggested by Scetbon et al. (2022), to promote the solver’s stability, we always initialize it with $\mathbf{P}_{\textrm{init}}=\tfrac{1}{n^{2}}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$.
@@
For usual costs, including inner products, $\ell_{p}^{q}$ distances, and standard CPD kernels, the computation of $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})$ (resp. $c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})$) can be done in $\mathcal{O}(d_{\mathcal{X}})$ time (resp. $\mathcal{O}(d_{\mathcal{Y}})$). Consequently, the DST can be computed in $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$ time. Furthermore, as discussed in § 2.2, the time complexity of the entropic GW solver is $\mathcal{O}(n^{3})$ in general, but can be reduced to $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$ when $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$, or $c_{\mathcal{X}}=\|\cdot-\cdot\|_{2}^{2},c_{\mathcal{Y}}=\alpha\|\cdot-\cdot\|_{2}^{2}$. Therefore, since the cosine similarity is equivalent to the inner product, up to pre-normalization of $\mathbf{x}_{i}$ and $T(\mathbf{x}_{i})$, this solver runs in $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$ for the costs of interest $\mathbf{(Sc)L2^{2}}$ and $\mathbf{Cos}$. The complete algorithm, along with a time complexity analysis of each step, is described in Alg. 1.
@@
[…]
@@ 3.5 Learning with the Gromov-Monge gap
[…]
@@ 4 Experiments
In Fig. 2, we start by illustrating the difference between the DST and GMG on synthetic data. We train $T_{\theta}$ to map from a $3$D source $p$ to a $2$D target distribution $q$ (first and second column) by minimizing $\mathcal{L}(\theta)\vcentcolon=\mathrm{S}_{\varepsilon}(T_{\theta}\sharp p,q)+\mathrm{R}(T_{\theta})$, where $\mathrm{S}_{\varepsilon}$ is the Sinkhorn divergence (Feydy et al., 2019) and $\mathrm{R}$ a geometry-preserving regularizer. We compare three settings, [i] no regularization $\mathrm{R}=0$, [ii] $\mathrm{R}=\mathrm{DST}_{p}$ as regularizer , and [iii] $\mathrm{R}=\mathrm{GMG}_{p}$. For all cases, we plot the transported distribution $T_{\theta}\sharp p$ after training. Without regularization (third column), we fit the marginal constraint $T_{\theta}\sharp p=q$ but do not preserve the geometric features. With the DST (fourth column), we preserve geometric features but do not fit the marginal constraint, i.e., $T_{\theta}\sharp p\neq q$. On the other hand, with the GMG (fifth column), we get the best compromise by approximating a Gromov-Monge map: we fit $T_{\theta}\sharp p=q$, while preserving the geometric features as fully as possible.
@@ 4.2 Leveraging the Gromov-Monge Gap for Disentanglement
Experimental Setup.  Having demonstrated how the DST enables (i) fitting a marginal constraint while (ii) preserving geometric features as much as possible, we now apply it to disentangled representation learning. Our primary goal is to investigate whether the GMG results in enhanced disentanglement compared to the DST by efficiently combining (i) prior matching with (ii) geometric constraints on the representation space. Moreover, we aim to determine which costs $c_{\mathcal{X}},c_{\mathcal{Y}}$ to choose in the geometric regularization and what part of the pipeline should be regularized, the encoder $e_{\phi}$, or the decoder $d_{\theta}$. The experimental setup is described below.
@@
  • Datasets. We benchmark over four standard datasets: Shapes3D (Kim and Mnih, 2018), DSprites Higgins et al. (2017), SmallNORB LeCun et al. (2004), and Cars3D Reed et al. (2015).
[…]
@@
+ Dec-GMG 70.0 ±$\pm 5.9$ 81.0 ±$\pm 3.2$ 93.3 ±$\pm 8.6$ 96.1 ±$\pm 3.8$
[…]
@@
DSprites Higgins et al. (2017)
Base $26.2$$\pm 18.5$ ${32.3}$$\pm 19.3$ $33.6$$\pm 17.9$ $48.7$$\pm 10.2$
@@
+ Dec-GMG 30.1 ±$\pm 5.6$ 36.4 ±$\pm 5.7$ 30.1 ±$\pm 5.6$ 36.4 ±$\pm 5.7$
GMG Consistently Enhances Disentanglement.  To further validate our findings, we benchmark the GMG for decoder regularization with angle preservation ($\mathbf{Cos}$) against its distortion counterpart across three more datasets. We report full results in Table 2. Again we observe that the GMG outperforms or performs equally well to its distortion equivalent, confirming the benefits of accounting for the optimal possible mapping in the regularization. Note that for SmallNORB and Cars3D, we found no benefits with respect to DCI-D in adding an HFS regularization and obtained the best results without it. We emphasize that using the GMG with $\mathbf{Cos}$ regularization significantly improves results for all datasets versus not using any isometric regularization. This establishes the GMG as an effective tool for enhanced disentanglement.
@@
Stability of the GMG. Furthermore, we investigate the stability of the GMG compared to the DST. Computing the GMG using Algorithm 1 involves solving an optimization problem through a GW solver. This experiment aims to demonstrate that our proposed method for solving the GW problem leads to a stable loss in the GMG . We assess the stability by measuring the alignment of GMG gradients $\nabla_{\theta}\mathrm{GMG}_{r_{n}}(T_{\theta})$ across 5 randomly sampled batches from each dataset $r_{n}\sim\mathcal{D}$, for a fixed neural map $T_{\theta}$. We repeat this procedure for each of the four datasets $\mathcal{D}$ considered and apply the same methodology to the DST. Figure 4 presents results for our best-performing configuration, which regularizes the decoder (i.e., $T_{\theta}=d_{\theta}$) and employs the $\mathbf{Cos}$ cost. Results for encoder regularization (i.e., $T_{\theta}=e_{\phi}$) are reported in Appendix D.1. We observe that the GMG’s gradients exhibit significantly higher alignment compared to those of the DST, demonstrating greater stability of our proposed regularizer. This suggests that the GMG ’s formulation, which accounts for the the minimal distortion, effectively mitigates the inherent variability inherent of the distortion, leading to more stable gradient computations.
@@ Acknowledgements
Co-funded by the European Union (ERC, DeepCell - 101054957). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Luca Eyring and Karsten Roth thank the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program for support. Karsten Roth also thanks the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for support. Zeynep Akata was supported by BMBF FKZ: 01IS18039A, by the ERC (853489 - DEXIM), by EXC number 2064/1 – project number 390727645. Fabian J. Theis consults for Immunai Inc., Singularity Bio B.V., CytoReason Ltd, Cellarity, and has ownership interest in Dermagnostix GmbH and Cellarity.

### v3

@@ 3.1 An Efficient Gap Formulation for Distortion
Recently, Uscidda and Cuturi (2023) introduced the Monge gap, a regularizer that measures whether a map $T$ transports a reference distribution at the minimal displacement cost. Building on this concept, we replace ”displacement” with ”distortion” to introduce the Gromov-Monge gap, a regularizer that assesses whether a map $T$ transports a reference distribution at the minimal distortion cost.
@@
The Influence of the Reference Distribution. A crucial property of $\mathrm{DST}_{r}$ is that if $T$ transforms $r$ without distortion, it will also apply distortion-free to any distribution $s$ whose support is contained within that of $r$. Formally, if $\mathrm{DST}_{r}(T)=0$ and $s\in\mathcal{P}(\mathcal{X})$ with $\mathrm{Spt}(s)\subseteq\mathrm{Spt}(r)$, then $\mathrm{DST}_{s}(T)=0$. This raises a question for the GMG: If $T$ maps $r$ with minimal distortion, does it similarly map $s$ with minimal distortion? We answer this question positively with Prop. (3.2).
@@ 3.2 Estimation and Computation from Samples
[…]
@@
1:Require: $\mathbf{x}_{1},\dots,\mathbf{x}_{n}\sim r$; map $T$; entropic solver GW, entropic regularization scale $\varepsilon_{0}$ (default =0.1=0.1), statistic operator on cost matrix stat (default $\texttt{mean}$).
[…]
@@
Choosing $\varepsilon=0$, we recover $\mathrm{GMG}_{{r}_{n},0}(T)=\mathrm{GMG}_{{r}_{n}}(T)$. Moreover, the entropic estimator preserves positivity, as for $\varepsilon\geq 0$, $\mathrm{GMG}_{r_{n},\varepsilon}(T)\geq 0$ (see B.3). We compute $\mathrm{GW}_{\varepsilon}({r}_{n},T\sharp{r}_{n})$ using Peyré et al. (2016)’s solver introduced in 2.2. We use the implementation provided by ott-jax (Cuturi et al., 2022). In practice, we select $\varepsilon$ based on (positive) statistics from the cost matrices $\mathbf{C}_{\mathcal{X}}$$\mathbf{C}_{\mathcal{Y}}$. We define a scale $\varepsilon_{0}$ and set $\varepsilon=\varepsilon_{0}\cdot\texttt{stat}(\mathbf{C}_{\mathcal{X}})\cdot\texttt{stat}(\mathbf{C}_{\mathcal{Y}})$. Standard options for the statistic include $\texttt{stat}\in\{\texttt{mean},\texttt{max},\texttt{std}\}$. This procedure is equivalent to running the entropic GW solver on the re-scaled cost matrices $\mathbf{C}_{\mathcal{X}}/\texttt{stat}(\mathbf{C}_{\mathcal{X}})$ and $\mathbf{C}_{\mathcal{Y}}/\texttt{stat}(\mathbf{C}_{\mathcal{Y}})$ with $\varepsilon=\varepsilon_{0}$; see App. B.2. It is a common practical trick, initially suggested for stabilizing the Sinkhorn algorithm Cuturi (2013).
@@
For usual costs, including inner products, $\ell_{p}^{q}$ distances, and standard CPD kernels, the computation of $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})$ (resp. $c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})$) can be done in $\mathcal{O}(d_{\mathcal{X}})$ time (resp. $\mathcal{O}(d_{\mathcal{Y}})$). Consequently, the DST can be computed in $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$ time. Furthermore, as discussed in § 2.2, the time complexity of the entropic GW solver is $\mathcal{O}(n^{3})$ in general, but can be reduced to $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$ when $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$, or $c_{\mathcal{X}}=\|\cdot-\cdot\|_{2}^{2},c_{\mathcal{Y}}=\alpha\|\cdot-\cdot\|_{2}^{2}$. Therefore, since the cosine similarity is equivalent to the inner product, up to pre-normalization of $\mathbf{x}_{i}$ and $T(\mathbf{x}_{i})$, this solver runs in $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$ for the costs of interest $\mathbf{(Sc)L2^{2}}$ and $\mathbf{Cos}$. The complete algorithm, along with a time complexity analysis of each step, is described in Alg. 1. We stress that, in all cases and for any cost function, the time complexity depends linearly on the dimensions of the source and target spaces, $d_{\mathcal{X}}$ and $d_{\mathcal{Y}}$, making the GMG scalable to high-dimensional distributions.
@@
As laid out, the GMG can be used as a regularization loss to push any model $T$ to move distributions with minimal distortion. A natural question arises: is this regularizer convex? In the following, we study the convexity of $T\mapsto\mathrm{GMG}_{r}(T)$ and its finite-sample counterpart $T\mapsto\mathrm{GMG}_{{r}_{n}}(T)$. We focus on the costs $\mathbf{L2}$ and $\mathbf{Cos}$. For simplicity, we replace $\mathbf{Cos}$ with $\langle\cdot,\cdot\rangle$, as these costs are equivalent, up to normalization of $r$ and $T$. We respectively denote by $\mathrm{GMG}_{r}^{2}$ and $\mathrm{GMG}_{r}^{\langle\cdot,\cdot\rangle}$ the GMG for these costs. We start by introducing a weaker notion of convexity, previously defined on $\mathbb{R}^{d}$ (Davis et al., 2018), which we extend here to $L_{2}(r)\vcentcolon=\{T\,|\,\|T\|_{L_{2}(r)}^{2}\vcentcolon=\int_{\mathcal{X}}\|T(\mathbf{x})\|_{2}^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})<+\infty\}$.
@@ 3.4 Learning with the Gromov-Monge gap
Figure 2: Learning of geometry-preserving maps with the DST and the GMG. Provided a mixture of Gaussian source distribution $p$ and a uniform or circular target distribution $q$, we minimize a fitting loss together with a geometry-preserving regularization. Each line correspond to a different data setup and cost. As costs on the top line, we use $\mathbf{L2^{2}}=\|\cdot-\cdot\|_{2}^{2}$, to preserve distances between points while on the bottom line, we use $\mathbf{Cos}=\textrm{cos-sim}(\cdot,\cdot)$, to preserve angles.
@@ 4 Experiments
In Fig. 2, we start by illustrating the difference between the DST and GMG. We train $T_{\theta}$ to map from a 33D source $p$ to a 22D target $q$ (first and second column) by minimizing $\mathcal{L}(\theta)\vcentcolon=\mathrm{S}_{\varepsilon}(T_{\theta}\sharp p,q)+\mathrm{R}(T_{\theta})$, where $\mathrm{S}_{\varepsilon}$ is the Sinkhorn divergence (Feydy et al., 2019). We compare three settings, [i] no regularization $\mathrm{R}=0$, [ii] $\mathrm{R}=\mathrm{DST}_{p}$ as regularizer , and [iii] $\mathrm{R}=\mathrm{GMG}_{p}$. For all cases, we plot the transported distribution $T_{\theta}\sharp p$ after training. Without regularization (third column), we fit the marginal constraint $T_{\theta}\sharp p=q$ but do not preserve the geometric features. With the DST (fourth column), we preserve geometric features but do not fit the marginal constraint, i.e., $T_{\theta}\sharp p\neq q$. On the other hand, with the GMG (fifth column), we get the best compromise by approximating a Gromov-Monge map: we fit $T_{\theta}\sharp p=q$, while preserving the geometric features as fully as possible.
@@ 4.2 Leveraging the Gromov-Monge Gap for Disentanglement
Experimental Setup.  Having demonstrated how the GMG enables (i) fitting a marginal constraint while (ii) preserving geometric features as much as possible, we now apply it to disentangled representation learning. Our primary goal is to investigate whether the GMG results in enhanced disentanglement compared to the DST by efficiently combining (i) prior matching with (ii) geometric constraints on the representation space. Moreover, we aim to determine which $c_{\mathcal{X}},c_{\mathcal{Y}}$ to choose and what part of the pipeline should be regularized, the encoder $e_{\phi}$, or the decoder $d_{\theta}$.
@@
  • Datasets. We benchmark over four $64\times 64$ image datasets: Shapes3D (Kim and Mnih, 2018), DSprites (Higgins et al., 2017), SmallNORB (LeCun et al., 2004), and Cars3D (Reed et al., 2015).
  • Hyperparameters. To ensure a fair experimental comparison, we follow recent works (Locatello et al., 2019a; 2020, Roth et al., 2023) by using the same architecture and hyperparameters. We perform a similar small grid search for the weighting terms: $\beta$ for the KL loss and $\gamma$ for HFS. Additionally, we include the weighting $\lambda$ for the geometry-preserving regularizer in the grid search. Note that we search over the same loss weightings $\lambda$ for DST, GMG, and Jac. For all experiments with the GMG, we compute it with Alg. 1, and systematically use $\varepsilon_{0}=0.1$ and $\texttt{stat}=\texttt{mean}$. As a result, $\varepsilon_{0}$ is not included in the grid search. We conduct an ablation study on $\varepsilon_{0}$ in App. D.3. We use a batch size of $n=64$. At this scale, the computational cost of compute the GMG loss for a batch is negligible, about 3 milliseconds. See App. C for full details on hyperparameters.
@@
+ Dec-GMG 70.0 $\pm 5.9$ 81.0 $\pm 3.2$ 93.3 $\pm 8.6$ 96.1 $\pm 3.8$
$\mathbf{Cos}:c_{\mathcal{X}}=c_{\mathcal{Y}}=\textrm{cos-sim}(\cdot,\cdot)$
@@
DSprites (Higgins et al., 2017)
Base $26.2$ $\pm 18.5$ 32.3{32.3} $\pm 19.3$ $33.6$ $\pm 17.9$ $48.7$ $\pm 10.2$
@@
+ Dec-GMG 30.1 $\pm 5.6$ 36.4 $\pm 5.7$ 30.1 $\pm 5.6$ 36.4 $\pm 5.7$
GMG Consistently Enhances Disentanglement.  To further validate our findings, we benchmark the GMG for decoder regularization with angle preservation ($\mathbf{Cos}$) against its distortion counterpart across three more datasets. We report full results in Table 2. Again we observe that the GMG outperforms or performs equally well to its distortion equivalent. Note that for SmallNORB and Cars3D, we found no benefits with respect to DCI-D in adding HFS and obtained the best results without it. We emphasize that using the GMG with $\mathbf{Cos}$ regularization significantly improves results for all datasets. This establishes the GMG as an effective tool for enhanced disentanglement.
@@
Stability of the GMG. Furthermore, we investigate the stability of the GMG compared to the DST. Computing the GMG using Algorithm 1 involves solving an optimization problem through a GW solver. This experiment aims to demonstrate that our proposed method for solving the GW problem leads to a stable loss in the GMG . We assess the stability by measuring the alignment of GMG gradients $\nabla_{\theta}\mathrm{GMG}_{r_{n}}(T_{\theta})$ across 5 randomly sampled batches from each dataset $r_{n}\sim\mathcal{D}$, for a fixed neural map $T_{\theta}$. We repeat this procedure for each of the four datasets $\mathcal{D}$ considered and apply the same methodology to the DST. Figure 4 presents results. We observe that the GMG’s gradients exhibit significantly higher alignment compared to those of the DST, demonstrating greater stability of our proposed regularizer. This suggests that the GMG, which accounts for the the minimal distortion, effectively mitigates the inherent variability inherent of the distortion, leading to more stable gradient computations.
@@ Acknowledgements
Co-funded by the European Union (ERC, DeepCell - 101054957). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Luca Eyring and Karsten Roth thank the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program for support. Karsten Roth also thanks the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for support. Zeynep Akata was supported by BMBF FKZ: 01IS18039A, by the ERC (853489 - DEXIM), by EXC number 2064/1 – project number 390727645. Fabian J. Theis consults for Immunai Inc., Singularity Bio B.V., CytoReason Ltd, Cellarity, and has ownership interest in Dermagnostix GmbH and Cellarity